# Conformal prediction for causal effects of continuous treatments

### v3

@@ D.1 Proofs of the supporting lemmas
We first prove that the problem (PS) fulfills the linear independence constraint qualifications. For all $i=m+1,\ldots,n+1$, we denote the constraints of problem (PS) as
@@
As a result, the constraints are linearly independent. This property suffices for the KKT conditions to hold at any (local) optimum of (PS).To furthermore show that the KKT conditions are also sufficient for a global optimum, we show that (PS) is Type-I invex. An optimization problem with objective function $f(x)$ and constraints $g(x)<=0$ with $x\in\mathbb{R}^{n+1}$ is Type-I invex at $x_{0}$, if there exists $\nu(x,x_{0})\in\mathbb{R}^{n+1}$, such that
@@
[20]. In problem (PS), the gradients of the objective function and of each constraint $h_{i}$ for all $i,j=m+1,\ldots,n+1$ at $x_{0}$ are given by
@@
the definition of Type-I invexity holds for (PS). Thus, the KKT conditions are also sufficient for a global optimum.
@@
We prove Theorem 4.2 in three steps: (i) We show that function class :=$\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$ indeed satisfies Eq. (2) for the intervention $a^{\ast}=a+\Delta_{A}$ and rewrite Eq. (5) as a convex optimization problem. (ii) We retrieve the corresponding dual problem, derive a dual prediction set, and show the equality of the coverage guarantee of the dual and the primal prediction sets. (iii) We derive $S^{\ast}$ from the dual prediction set to construct $C_{n+1}$ and prove the overall coverage guarantee. For further theoretical background on the idea of the proof, we refer to Gibbs et al. [18].
@@
Observe that $\mathbb{E}[f(A,X)]=\theta$ for all $f\in\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$. Therefore, Eq. (2) simplifies to
@@
Thus, $\mathcal{F}$ satisfies the propensity shift from Eq. (2) for the soft intervention $a^{\ast}=a+\Delta_{A}$. Following Lemma 4.1, we thus aim to find
@@
Setting derivative of $\mathcal{L}$ w.r.t. $u_{i}$ and $v_{i}$ to 00 results in
@@
=$\quad$= $\tilde{S}u^{S}_{n+1}-Su^{\tilde{S}}_{n+1}+\tilde{S}g^{S}(x_{n+1},a_{n+1})-Sg^{\tilde{S}}(x_{n+1},a_{n+1})+\tilde{S}v^{\tilde{S}}_{n+1}-Sv^{S}_{n+1}$ (102)
<$\quad$< $S(u^{S}_{n+1}-u^{\tilde{S}}_{n+1}+g^{S}(x_{n+1},a_{n+1})-g^{\tilde{S}}(x_{n+1},a_{n+1})-(v^{S}_{n+1}-v^{\tilde{S}}_{n+1}))$ (103)
=$\quad$= $S(S-\tilde{S}).$ (104)
@@ E.1 Extended literature review
Conformal prediction under covariate shift Multiple works on CP with marginal coverage under distribution shifts between training and test data have been introduced in the literature [e.g., 8, 11, 15, 16, 17, 18, 19, 36, 42, 50, 57]. Our setting also involves a distribution shift due to the intervention on the treatment but differs from the latter in that the true distribution shift is unknown.
@@
E.2 Discussion on the tightness of our CP intervals:
@@
E.3 A note on challenges and difficulties in CP for causal effects of continuous treatments
@@
E.4 Causal effects of continuous treatments & kernel smoothing
@@
E.5 Interpretation of optimal parameters
@@
E.6 Interpretation of parameter $M$
Our optimization requires the specification of a parameter $M$, denoting a bound on the propensity estimation error. One can view the parameter $M$ as a type of sensitivity parameter. Therefore, we follow former work in causal inference and propose to incorporate domain knowledge to specify the parameter $M$ [e.g., 13, 48]. Another way of making use of the parameter $M$ is to observe how the intervals change for varying $M$. This indicates how much effect the propensity misspecification has on the prediction interval and can help in making reliable decisions. A third option to calibrate $M$ is to employ measures for epistemic uncertainty on top of the propensity estimate when there is no domain knowledge for specifying $M$.
E.7 A note on the stability of our method
@@
Appendix F Experimentation details
F.1 Synthetic dataset generation
@@ F.3 Implementation details
F.4 Selection of the interventions in our experiments
@@
[…]
Appendix G Further results
@@
We further report the width of the prediction intervals in our synthetic experiments in Table 5. The width is important to assess the usefulness of the resulting prediction intervals. As the performance of the ensemble method is not comparable with the coverage of MC-Dropout and our CP method, we only compare the latter two methods with regard to the interval width.
@@
Scalability: Calculating the prediction intervals requires an iterative search for an optimal value $S^{\ast}$. Therefore, the underlying optimization problem must be fitted multiple times throughout the algorithm, potentially posing scalability problems. In our empirical studies, however, we did not encounter scalability issues. Importantly, we found that the average runtime of our algorithm on a standard desktop CPU is only 16.43 seconds.

### v4

@@ D.1 Proofs of the supporting lemmas
We first prove that the problem equation 18 fulfills the linear independence constraint qualifications. For all $i=m+1,\ldots,n+1$, we denote the constraints of problem equation 18 as
@@
As a result, the constraints are linearly independent. This property suffices for the KKT conditions to hold at any (local) optimum of equation 18.To furthermore show that the KKT conditions are also sufficient for a global optimum, we show that equation 18 is Type-I invex. An optimization problem with objective function $f(x)$ and constraints $g(x)<=0$ with $x\in\mathbb{R}^{n+1}$ is Type-I invex at $x_{0}$, if there exists $\nu(x,x_{0})\in\mathbb{R}^{n+1}$, such that
@@
[20]. In problem equation 18, the gradients of the objective function and of each constraint $h_{i}$ for all $i,j=m+1,\ldots,n+1$ at $x_{0}$ are given by
@@
the definition of Type-I invexity holds for equation 18. Thus, the KKT conditions are also sufficient for a global optimum.
@@
We prove Theorem 4.2 in three steps: (i) We show that function class $\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$ indeed satisfies Eq. equation 2 for the intervention $a^{\ast}=a+\Delta_{A}$ and rewrite Eq. equation 5 as a convex optimization problem. (ii) We retrieve the corresponding dual problem, derive a dual prediction set, and show the equality of the coverage guarantee of the dual and the primal prediction sets. (iii) We derive $S^{\ast}$ from the dual prediction set to construct $C_{n+1}$ and prove the overall coverage guarantee. For further theoretical background on the idea of the proof, we refer to Gibbs et al. [18].
@@
Observe that $\mathbb{E}[f(A,X)]=\theta$ for all $f\in\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$. Therefore, Eq. equation 2 simplifies to
@@
Thus, $\mathcal{F}$ satisfies the propensity shift from Eq. equation 2 for the soft intervention $a^{\ast}=a+\Delta_{A}$. Following Lemma 4.1, we thus aim to find
@@
Setting derivative of $\mathcal{L}$ w.r.t. $u_{i}$ and $v_{i}$ to 0 results in
@@
$=\quad$ $\tilde{S}u^{S}_{n+1}-Su^{\tilde{S}}_{n+1}+\tilde{S}g^{S}(x_{n+1},a_{n+1})-Sg^{\tilde{S}}(x_{n+1},a_{n+1})+\tilde{S}v^{\tilde{S}}_{n+1}-Sv^{S}_{n+1}$ (103)
$<\quad$ $)\displaystyle S(u^{S}_{n+1}-u^{\tilde{S}}_{n+1}+g^{S}(x_{n+1},a_{n+1})-g^{\tilde{S}}(x_{n+1},a_{n+1})-(v^{S}_{n+1}-v^{\tilde{S}}_{n+1}$)) (104)
[…]
@@ E.1 Extended literature review
Conformal prediction under covariate shift Multiple works on CP with marginal coverage under distribution shifts between training and test data have been introduced in the literature [e.g., 8, 11, 15, 16, 17, 18, 19, 36, 42, 49, 56]. Our setting also involves a distribution shift due to the intervention on the treatment but differs from the latter in that the true distribution shift is unknown.
@@
E.2 The need for exchangeability in CP
Coverage guarantees of existing CP intervals essentially rely on the exchangeability of the non-conformity scores. Exchangeability assures that the nonconformity score of the test point $n+1$ is equally likely to fall anywhere among the calibration scores, its rank is uniform, and that uniformity is exactly what yields the distribution‑free coverage guarantee. Without exchangeability, the rank is not guaranteed to be uniform, and the marginal coverage bound can fail.
However, intervening on treatment $A$ shifts the propensity function and, therefore, induces a shift in the covariates between calibration and test data, specifically in treatment $A$. Therefore, exchangeability is not fulfilled, and the coverage guarantees might fail.
As a remedy, we present a novel and powerful remedy in our work: The overall distribution of the confounders $X$ is assumed to stay constant between train, calibration, and test data (as standard in ML problems). This is completely orthogonal to constructing intervals for different (e.g., young or old) patients. Note that CP intervals are constructed for only one sample/patient at a time. This means that different intervals are constructed for patients with different features $X$. In other words, the intervals are constructed conditionally on $X$, but the coverage guarantee is marginal across the complete population of $X$. Overall, the shift from one patient to another does not pose any challenges for CP methods.
Appendix F Extended discussion
F.1 Discussion on the tightness of our CP intervals:
@@
F.2 A note on challenges and difficulties in CP for causal effects of continuous treatments
@@
F.3 Causal effects of continuous treatments & kernel smoothing
@@
F.4 Interpretation of optimal parameters
@@
F.5 Interpretation of parameter $M$
Our optimization requires the specification of a parameter $M$, denoting a bound on the propensity estimation error. One can view the parameter $M$ as a type of sensitivity parameter. Therefore, we follow former work in causal inference and propose to incorporate domain knowledge to specify the parameter $M$ [e.g., 13, Tan.2006]. Another way of making use of the parameter $M$ is to observe how the intervals change for varying $M$. This indicates how much effect the propensity misspecification has on the prediction interval and can help in making reliable decisions. A third option to calibrate $M$ is to employ measures for epistemic uncertainty on top of the propensity estimate when there is no domain knowledge for specifying $M$.
F.6 A note on the stability of our method
@@
Appendix G Experimentation details
G.1 Synthetic dataset generation
@@ G.3 Implementation details
G.4 Selection of the interventions in our experiments
@@
We sampled the covariates $X$ in the range from 1 to 4. To again perform interventions that fall inside the range of [0,40], we decided to set the intervention as 7XX and 10XX. Other choices of interventions would also have been possible. Reassuringly, there was no systematic selection of interventions in our experiments besides the considerations above.
For the soft interventions, we do not see different effects of the interventions on the two datasets. For the hard interventions, however, we can observe a difference. Recall that in dataset 1, the treatment was sampled uniformly from [0,40] (with a dependence on $X$), whereas in dataset 2, it was sampled from a normal distribution with a mean of 5XX. Therefore, the intervention 10XX is far in the tail of the distribution. As a result, we observe a slightly lower coverage for this intervention on dataset 2 compared to dataset 1.
Appendix H Further results
@@
Width of the prediction intervals: We further report the width of the prediction intervals in our synthetic experiments in Table 5. The width is important to assess the usefulness of the resulting prediction intervals. As the performance of the ensemble method is not comparable with the coverage of MC-Dropout and our CP method, we only compare the latter two methods with regard to the interval width.
@@
Comparison to the vanilla CP baseline: We compare our method to the naive vanilla CP (V-CP), i.e., a CP method that does not account for the distribution shift. We observe that V-CP does not achieve any valid prediction interval across all distribution shifts and confidence levels. This can be explained by the good prediction performance of the underlying model. Thus, V-CP intervals are extremely small (average width of 0.0003) and can never cover the true potential outcome after the intervention. Overall, the results confirm the importance of accounting for the distribution shift induced by the intervention.
@@
Scalability: Calculating the prediction intervals requires an iterative search for an optimal value $S^{\ast}$. Therefore, the underlying optimization problem must be fitted multiple times throughout the algorithm, potentially posing scalability problems. In our empirical studies, however, we did not encounter scalability issues. Importantly, we found that the average runtime of our algorithm on a standard desktop CPU is only 16.43 seconds. On the MIMIC dataset, computing CP intervals takes roughly 10 times longer than computing MC intervals. However, we emphasize that MC intervals are generally not faithful and therefore not directly comparable.